Carolina Company uses the LIFO method for valuing its ending inventory. The following financial statement information is available for its first year of operation: Carolina Company Income Statement For the year ended December 31 Sales 60,000 Cost of Goods sold 23,000 Gross Profit 37,000 Expenses 13,000 Income before taxes $ 24,000 Carolina's ending inventory using the LIFO method was $8,700. Carolina's accountant determined that had the company used FIFO, the ending inventory would have been $9,100

. a. Determine what the income before taxes would have been, had Carolina used the FIFO method of inventory valuation instead of LIFO. b. What would be the difference in income taxes between LIFO and FIFO, assuming a 30% tax rate? c. If Carolina wanted to lower the amount of income taxes to be paid, which method would it choose?
Answer #1

a) Cost of goods available for sale= Cost of goods sold+Ending inventory

Under LIFO method

Cost of goods available for sale= $23000+8700= $31700

Under FIFO method

Cost of goods available for sale= Cost of goods sold+Ending inventory

$31700= Cost of goods sold+9100

Cost of goods sold= 22600

So, income before tax under FIFO method

Income before tax= Sales-Cost of goods sold-Expenses

= $60000-22600-13000= $24400

b) Tax= Income before taxes*Tax rate

Under LIFO method, tax= $24000*30%= $7200

Under FIFO method, tax= $24400*30%= $7320

Difference in income taxes between LIFO and FIFO method= $7320-7200= $120

c) LIFO method should be selected, as the Carolina wanted to lower the amount of income tax to be paid, as the income before taxes under LIFO method is lower than the income before FIFO method that means the amount of income taxes to be paid under LIFO method will be lower than FIFO method.